OpenStudy (anonymous):

Find all points on the graph of f(x)=-x^3 + 3x^2 - 2 at which there is a horizontal tangent line.

OpenStudy (anonymous):

I feel bad to have skipped this one earlier. Anyway, you must first find the first derivative. f'(x) = -3x^2 + 6x Then you must find the points at which f'(x) = 0. f'(x) = 3x(3-x) 0 = 3x; x = 0 0 = 3-x; x = 3 x = 0 and 3

OpenStudy (anonymous):

Okay, but where did you get 3x(3-x)? And is it okay if I ask you another question in this question? I'm not sure if that's allowed or not, I just started using this site haha.

OpenStudy (anonymous):

I factored out a 3x from the derivative